(a) A company produces and sells television sets at RM 350 each. The company’s fixed costs
per annum are as follows:
RM
Rent 3,040
Heating and lighting 2,760
Salaries 14,000
Advertising 900

The variable costs are RM 120 per television set. Using calculation method, find;

(i) the break-even point, (4 marks)

(ii) the loss when output is 70, (3 marks)

(iii) the profit when output is 300. (3 marks)
1
Expert's answer
2015-09-16T11:44:49-0400
(a) P = 350, FC = 3,040 + 2,760 + 14,000 + 900 = 20,700,
AVC = 120.
(i) the break-even point will be Q = FC/(P - AVC) = 20,700/(350 - 120) = 20,700/230 = 90 units.
(ii) the loss when output is 70 will be TP = TR - TC = P*Q - (FC + AVC*Q) = 350*70 - (20,700 + 120*70) = -4,600.
(iii) the profit when output is 300 will be TP = TR - TC = P*Q - (FC + AVC*Q) = 350*300 - (20,700 + 120*300) = 48,300.